OpenStudy (anonymous):

Which expression represents how many years are in x months? A) 12x B) 12 + x C) x 12 D) 12 x

OpenStudy (anonymous):

im guessing a not sure

OpenStudy (anonymous):

There should be something else with this problem

OpenStudy (anonymous):

there strangly wasn't

OpenStudy (anonymous):

@arabpride ^_^

OpenStudy (tylerd):

1 year = 12 months 1 month = 1year/12

OpenStudy (tylerd):

so if months = x 1 year = 12x x = 1/12 years

OpenStudy (anonymous):

A) 12x B) 12 + x C) x/12 D) 12/x

OpenStudy (anonymous):

intrestingg..

OpenStudy (anonymous):

@skittykitty6969

OpenStudy (tylerd):

what do you think it is

OpenStudy (anonymous):

ummm A.

OpenStudy (tylerd):

try x/12

OpenStudy (tylerd):

if we say 6 months we have 6/12 = 1/2 a year

OpenStudy (tylerd):

theres 1/2 a year in 6 months so its true

OpenStudy (anonymous):

ohhh man your right oppps thank you!!!:D

OpenStudy (anonymous):

oh crap im sorry i went out to eat